OpenStudy (anonymous):

Which equation could be used to solve the problem? Joan picked 4 bushels of apples in one hour. At that rate, how many bushels would she pick in 8 hours (b)?

OpenStudy (anonymous):

A. b = 8 + 4 B. b = 8 ÷ 4 C. b = 8 • 4 D. b = 4 ÷ 8

OpenStudy (anonymous):

is it b

OpenStudy (anonymous):

why you think so

OpenStudy (anonymous):

because i know its divison

OpenStudy (anonymous):

how you know that, who wispered it to you? Try to think about it for a momento, and try again

OpenStudy (anonymous):

can u just tell me idk

OpenStudy (anonymous):

or is it c

OpenStudy (anonymous):

this question is not really about math, it is just a común sence. I am sure you can answer it yourself if you just give it a try

OpenStudy (anonymous):

ok so its not b

OpenStudy (anonymous):

you are like in tv show, tring to guess under which hat the ball is hidden

OpenStudy (anonymous):

my finel gusses is c

OpenStudy (anonymous):

this is not a show, and it is not about guessing

OpenStudy (anonymous):

can u just tell me im getting upset

OpenStudy (anonymous):

i have looked at this like 20 times plzs

OpenStudy (anonymous):

think of it for a momment. Immagine yourself picking those apples. 1 hour passed and you picked up 4 bushels. If you keep doing for another hour, how many more will you pick up?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i have a nother question

OpenStudy (anonymous):

if you pick up 4 in 1 hour, and you it at the same speed for two hours, you eill pick up 8 bushels.

OpenStudy (anonymous):

A wheel made 45 revolutions in 1 minute. At that rate, how many revolutions would it make in 15 minutes?

OpenStudy (anonymous):

would i do 45*15=675

OpenStudy (anonymous):

so: 1 hour - 4 bushels 2 hours - 8 bushels 3hours - 12 bushels 4 hours - 16 bushels and so on

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the answer is 8*4=? and I really hope you try to think why it is so. C ya